OpenStudy (anonymous):

will the expression 5/x always be less than5 explain

myininaya (myininaya):

false: counterexample: x=5^-2 so we 5/(5^-2)=5(5)^2=125>5

OpenStudy (anonymous):

i am in the 7th grade i don't understand your repy, plz help

myininaya (myininaya):

do you know law of exponents?

OpenStudy (anonymous):

no

myininaya (myininaya):

ok what happens of x=0.1 what is 5/x=____

myininaya (myininaya):

if*

myininaya (myininaya):

what is 5/(0.1)=____

OpenStudy (anonymous):

what is 5 divided by 0.1?

myininaya (myininaya):

yes

myininaya (myininaya):

5/(0.1) you could multiply the top and bottom by 10 and if you did what is this fraction equivalent to____

myininaya (myininaya):

what is 5*10=____

OpenStudy (anonymous):

50

myininaya (myininaya):

what is 0.1*10=___

OpenStudy (anonymous):

.10

myininaya (myininaya):

no 0.1*10=1 0.01*10=0.1 0.001*10=0.01 what is 0.0001*10=____

OpenStudy (anonymous):

is that divided?

myininaya (myininaya):

* means multiply

OpenStudy (anonymous):

ok so the answer would be no because it can be broken down in a simpler form

myininaya (myininaya):

i don't know what you are talking about, but the question is 5/x always less than 5 is false. A counterexample would be for x=0.1 since 5/(0.1)=5/(0.1)*(10/10)=50/1=50 and 50>5

myininaya (myininaya):

there are many different counterexamples

myininaya (myininaya):

but you only need one

OpenStudy (anonymous):

ok so it can be many other examples

myininaya (myininaya):

there are some examples that will show this is true( so the always part is false since there are also examples showing that the statement is false)

OpenStudy (anonymous):

thank you is math your favorite subject?

myininaya (myininaya):

yes lol

OpenStudy (anonymous):

ok great thank you, i have tons of other questions to finish

myininaya (myininaya):

ok have fun

OpenStudy (anonymous):

ok